solve using the subtuition method 3x-2y=8
5x+3y=7 find the value of y

Answer by mananth(16946)

3 x + -2 y = 8 .............1
-2 y = 8 + -3 x
/ -2 =
y = -4 + 1.5 x
5 x + 3 y = 7 .....................2
Substitute y in (2)
5 x + 3 ( -4 + 1.5 x ) =
5 x + -12 + 4.5 x = 7
5 x 4.5 x = 19
9.5 x = 19
/ 9.5
x= 2
Plug the value of x in (1)
3 x + -2 y = 8
3 * 2 + -2 y = 8
6 + -2 y = 8
-2 y = 2
/ -2
y= -1
m.ananth@hotmail.ca